40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 253 of 344  Not logged in ELOG logo
ID Date Author Type Category Subject
  4595   Mon May 2 12:51:52 2011 kiwamuUpdateLSCPRMI locking : plan

Today we will try to lock the PRMI. Here is a plan for it.

(to be done in the daytime)

 - setup REFL11 RFPD

 - setup AS55 RFPD

 - install a demod board for 55 MHz

 - install a 3-way RF combiner on EOM.

 - prepare 55 MHz RF source (Marconi or RF source box ?)

 - adjustment of each demodulation phase

 - activation of PRM oplev

 

 (control topology)

- AS55_Q ==> BS

- REFL11_I ==> PRM

  4594   Mon May 2 11:14:27 2011 kiwamuUpdateSUSscript : opticshutdown

Just FYI.

There is a useful script for this particular job : shutting down all the suspensions and bringing it back to operation after 5 hrs.

It is called opticshudown, which resides in /cvs/cds/rtcds/caltech/c1/scripts/SUS/.

 

Also I added this script on the list in the wiki where all the scripts will be listed.

http://blue.ligo-wa.caltech.edu:8000/40m/Computers_and_Scripts/All_Scripts#opticshutdown

If you find any other useful scripts, please add them on the wiki.

Quote from #4592

I leave all the suspensions free from the watchdogs for 5 hours from now.

 

  4593   Sat Apr 30 05:14:33 2011 KojiUpdateGreen LockingElectronics schematic for vertex beatbox

- AC coupling for the comparator circuit of the green locking

In order to relieve the power consumption of the RF buffer, ac coupling circuits have been added.

The ac coupling before the buffer amp helps to relieve the power consumption in the chip.
But because of the distortion of the signal (and the limitation of the bandwidth), the output still has some DC (~0.6V).
Therefore, the output is also AC coupled.

Note that the BW pin of BUF634P should be directly connected to -15V in order to keep the bandwidth of the buffer.

The drawings are also uploaded on the green electronics wiki

Attachment 1: freq_divider.png
freq_divider.png
  4592   Sat Apr 30 00:15:26 2011 kiwamuUpdateSUSDRMI prep : suspension diagnostic

As a part of the DRMI preparation,

I leave all the suspensions free from the watchdogs for 5 hours from now.

Please DO NOT touch them.

I will check the spectra and the mechanical resonant frequencies on Monday.

Also I will renew all the input matrices of the local dampings based on these free swinging spectra.

  4591   Fri Apr 29 18:24:05 2011 SureshUpdateRF SystemRF system: 1X2 Rack cabling

[Joe, Jamie, Suresh]

We have installed the IDE to SCSI adaptor module into the 1X2 rack and have connected the AA filter outputs to it.

P4290070.JPG

 

We have removed the following cables running between the 1X2 and 1X3 racks.

The long twisted pair ribbon cable which previously carried the ADC signals.

1X2-ASC 6, 1X2-ASC 47, 1X2-ASC 9, 1X2-ASC 8, 1X2-ASC 10, 1X2-ASC 7,

CAB-1X2-LSC 42, CAB 1X2-LSC 56,  CAB 1X2-LSC 41, CAB 1X2-LSC 43

1X3-2 ASC 47

We have also removed the following by mistake.  We will put them back them on Monday

1X2-LSC 21, 1X2-LSC-20.

We have also removed the ASC QPD cables and moved the QPD cards which were present in the middle Eurocate (#2) to the unused Eurocrate at the bottom position (#3).

The binary input cables at the back of the cards require to be supported so that their weight does not pull them out of the sockets at the back of the crates.

Some of the slots where we plan to plug in Demod boards (the 165 MHz boards)  are not currently connected to any binary output on the C1:LSC computer.  We need these binary controls for the fitlter modules on the cards.

When we eventually begin to use the 15PDs as planned, then we will occupy 30 ADC channels (I & Q outputs).  Currently we have just one ADC card installed on the C1:LSC providing 32 ADC channels.  Joe found another 16bit 32 channel ADC card in his stash but we need to get a timing+adaptor board for it. In general we are going to need the third Eurocrate.

A platform for the power supply of the RF Distribution box needs to be built and the power supply needs to be moved into the 1X2 rack rather than sit on top of 1X2 rack.

 

 

 

  4590   Fri Apr 29 14:36:36 2011 josephb, steveUpdateCDSRemoved hanging D-sub to SCSI in 1X5

Quote:

Problem:

Way back, Jay had D-sub to SCSI adapters made to adapt our existing Sander box AA filters to the new SCSI based IO chassis.  However, these did not fit inside the box.

At the time, we simply left the cards outside hanging, which was a hack and needed to be replaced.

Solution:

Steve modified a black AA filter box so that it could fit the D-sub to SCSI adapter board on it, plus strain relief the SCSI cable, rather than let it hang.  The back of the box was cut, and an extending piece of metal attached to the bottom of the box.  The adapter board was screwed into the box, the SCSI plugged in, then the SCSI cable is clamped to the extending metal as well.

This modification will be propagated to the 3 remaining AA filter boards using the D-sub to SCSI adapter.

 The same modification was carried out at 1X5 for PRM & SRM.

Note:  D68L8EX-850Hz  are removed  and bypassed in 7 channels.

Attachment 1: P1070621.JPG
P1070621.JPG
  4589   Fri Apr 29 13:58:36 2011 ranaUpdateLSCY arm locked : details

Quote:

 * Farfalla, a lab laptop, seems out of network.

 If you look at the real host table instead of the misleading host table in the wiki, you will see that someone has deleted Farfalla from there. She needs to be re-added.

  4588   Fri Apr 29 13:15:04 2011 kiwamuUpdateLSCY arm locked : details

As far as I know, this button works only once after the launch of MEDM...

Quote:

 * reload button on sitemap.adl doesn't work. 

 

  4587   Fri Apr 29 12:18:48 2011 kiwamuUpdateLSCY arm locked : details

First of all, the conclusions / results from the exercise of the Y arm locking yesterday are:

   The position of the beam spots on both ETMY and ITMY are now not so bad ( ~ 5 mm off from the center).

   The input PZTs are coarsely aligned to the Y arm.

   Nevertheless IP_ANG is still too high to come out from the view port at the Y end station.

  After the alignments of PRM, SRM and Michelson, POP is still largely clipped.

 


(what I did)

  - Alignments of the Input PZTs

    First I tried letting the incident beam hit the center of ETMY by steering PZT1 and 2 as usual.
    Then I coarsely aligned the cavities to the incident beam and checked the beam flashing spots on ETMY and ITMY with the CCD monitors.
    When the spots were far from the center I went back to the alignment of PZT1 and 2 to get better beam positions. And repeated this work several times.

  - Adjustment of the demodulation phase  for the Y arm PDH.

    First I started looking at the digital signals and tried correcting the demodulation phase by the rotation matrix, but this didn't go fast because I had to do some DAQ settings, plotting and analysis.
    Instead looking at the digital signal, I observed the analog signals with an oscilloscope. I found the demodulation phase was something like 45 deg.
    Based on the analog measurement I rotated the digital matrix by 45 deg to get the I-signal maximized. Indeed this worked well. I obtained a beautiful PDH signal from the I-signal.
    Note that we are using 11 MHz mod/demod and eventually the signals come out from "REFL33" on the digital side.

  - Activation of oplev on ITMY

    Instead using ETMY I used ITMY for the length control because somehow I felt that ETMY coils were suspicious and they looked not so nice.
    One of the reason is that ETMY's coil actuation efficiencies looked low compared to the other test masses.
    For example a gain of 700 for SUSSIDE damping is needed on ETMY to get a reasonable Q. This is about 2 - 3 times larger gain than the other test masses.
    So I started using ITMY for the locking and activated the oplev to suppress unwanted excitation due to kicks from the control signal during the locking,
    The oplev has been misaligned, so I went to the ITMY optical bench and tweaked a steering mirror to let the He-Ne beam go into the QPD.
    I set the gains +2 for PITCH and -2 for YAW.

  - PDH locking

    The locking had been quite difficult even though the cavity alignment was quite good.
    It's because the beam on AS11_RFPD was almost falling off from the photo diode. This causes a big amplitude fluctuation in the PDH signal as the beam position moves.
    After aligning the beam by steering BS I got able to lock the Y arm. The PD whitening gains are all 0 dB and the feedback gain is -2, giving us a UGF of 250 Hz.

 

(Broken or likely broken stuff)

 * IP_ANG doesn't give a signal to the digital side.

 * ETMY coils look weak and 2 - 3 times weaker than the other test masses. (or OSEM readout gain maybe lower)

 * reload button on sitemap.adl doesn't work.

 * Farfalla, a lab laptop, seems out of network.

Quote from #4586

The Y arm has been locked with the IR beam. The purpose is to use the arm as an alignment reference for the input PZTs.

Detail will be posted later. Here is a picture of ITMY suspension. You can see there is a beam spot in the middle of the test mass.

 

  4586   Fri Apr 29 05:48:52 2011 kiwamuUpdateLSCY arm locked

The Y arm has been locked with the IR beam. The purpose is to use the arm as an alignment reference for the input PZTs.

Detail will be posted later. Here is a picture of ITMY suspension. You can see there is a beam spot in the middle of the test mass.

DSC_2978_ss.jpg

  4585   Fri Apr 29 03:39:49 2011 KojiSummaryLSCCavity lengths

I tried the idea that the PRC can resonate f1 and f2 at the same time if the arm gives the reflection phase to f1 and f2 with the ratio of 1 vs 5.

The details are described on wiki. The point is this removes all of the PRC/SRC/asymmetry mumbo jumbo.

The calculated cavity lengths for f_mod of 11.065399MHz are:

  • Arm Length: 37.7974 [m]

  • PRC Length: 6.7538 [m]

  • SRC Length: 5.39915 [m]

  • Asymmetry (lx-ly): 0.0342 [m]


Here is the actual values derived from the photos.

  • Arm Length: 37.54 [m] (0.26m too short)

  • PRC Length: 6.760 [m] (6mm too long)

  • SRC Length: 5.415 [m] (16mm too long)

  • Asymmetry (lx-ly): 0.0266 [m] (8mm too long)

  4584   Thu Apr 28 22:38:38 2011 AidanUpdateGreen LockingElectronics schematic for vertex beatbox

 With some assistance from Kiwamu and Koji, I've drawn up the electronics design for the Beat Box for the vertex green locking. The Omingraffle schematic is posted on the Green Locking Wiki page. It's also attached below. Some final touches are necessary before we can Altium this up.

 Attachment 1: Schematic of beatbox

Attachment 2: Front and back panel designs.

Attachment 1: Canvas_1.png
Canvas_1.png
Attachment 2: Canvas_2.png
Canvas_2.png
  4583   Thu Apr 28 16:12:19 2011 josephb, jamieUpdateCDSNew CDS model SVN

New SVN

We are now using the LIGO CDS SVN for storing our control models.

The SVN is at:

https://redoubt.ligo-wa.caltech.edu/websvn/

The models are under cds_user_apps, then trunk, then approriate subsystem (ISC for c1lsc for example), c1 (for caltech 40m), then models.

We have checked out /cds_user_apps to /opt/rtcds/.

So to find the c1lsc.mdl model, you would go to /opt/rtcds/cds_user_apps/trunk/ISC/c1/models/c1lsc.mdl

This SVN is shared by many people LIGO, so please follow good SVN practice.  Remember to update models ("svn update") before doing commits.  Also, after making changes please do an update to the SVN so we have a record of the changes.

New Practices

We are creating soft links in the /opt/rtcds/caltech/c1/core/advLigoRTS/src/epics/simLink/ to the models that you need to build.  So if you want to add a new model, please add it to the cds_users_apps SVN in the correct place and create a soft link to the simLink directory.

lrwxrwxrwx 1 controls controls      54 Apr 28 14:41 c1sus.mdl -> /opt/rtcds/cds_user_apps/trunk/SUS/c1/models/c1sus.mdl
lrwxrwxrwx 1 controls controls      54 Apr 28 14:41 c1sup.mdl -> /opt/rtcds/cds_user_apps/trunk/SUS/c1/models/c1sup.mdl
lrwxrwxrwx 1 controls controls      54 Apr 28 14:41 c1spy.mdl -> /opt/rtcds/cds_user_apps/trunk/SUS/c1/models/c1spy.mdl
lrwxrwxrwx 1 controls controls      54 Apr 28 14:41 c1spx.mdl -> /opt/rtcds/cds_user_apps/trunk/SUS/c1/models/c1spx.mdl
lrwxrwxrwx 1 controls controls      54 Apr 28 14:41 c1scy.mdl -> /opt/rtcds/cds_user_apps/trunk/SUS/c1/models/c1scy.mdl
lrwxrwxrwx 1 controls controls      54 Apr 28 14:41 c1scx.mdl -> /opt/rtcds/cds_user_apps/trunk/SUS/c1/models/c1scx.mdl
lrwxrwxrwx 1 controls controls      54 Apr 28 14:41 c1mcs.mdl -> /opt/rtcds/cds_user_apps/trunk/SUS/c1/models/c1mcs.mdl
lrwxrwxrwx 1 controls controls      54 Apr 28 14:41 c1x05.mdl -> /opt/rtcds/cds_user_apps/trunk/CDS/c1/models/c1x05.mdl
lrwxrwxrwx 1 controls controls      54 Apr 28 14:41 c1x04.mdl -> /opt/rtcds/cds_user_apps/trunk/CDS/c1/models/c1x04.mdl
lrwxrwxrwx 1 controls controls      54 Apr 28 14:41 c1x03.mdl -> /opt/rtcds/cds_user_apps/trunk/CDS/c1/models/c1x03.mdl
lrwxrwxrwx 1 controls controls      54 Apr 28 14:41 c1x02.mdl -> /opt/rtcds/cds_user_apps/trunk/CDS/c1/models/c1x02.mdl
lrwxrwxrwx 1 controls controls      54 Apr 28 14:41 c1x01.mdl -> /opt/rtcds/cds_user_apps/trunk/CDS/c1/models/c1x01.mdl
lrwxrwxrwx 1 controls controls      54 Apr 28 14:41 c1rfm.mdl -> /opt/rtcds/cds_user_apps/trunk/CDS/c1/models/c1rfm.mdl
lrwxrwxrwx 1 controls controls      55 Apr 28 14:41 c1dafi.mdl -> /opt/rtcds/cds_user_apps/trunk/CDS/c1/models/c1dafi.mdl
lrwxrwxrwx 1 controls controls      54 Apr 28 14:41 c1pem.mdl -> /opt/rtcds/cds_user_apps/trunk/ISC/c1/models/c1pem.mdl
lrwxrwxrwx 1 controls controls      54 Apr 28 14:41 c1mcp.mdl -> /opt/rtcds/cds_user_apps/trunk/ISC/c1/models/c1mcp.mdl
lrwxrwxrwx 1 controls controls      54 Apr 28 14:41 c1lsp.mdl -> /opt/rtcds/cds_user_apps/trunk/ISC/c1/models/c1lsp.mdl
lrwxrwxrwx 1 controls controls      54 Apr 28 14:41 c1lsc.mdl -> /opt/rtcds/cds_user_apps/trunk/ISC/c1/models/c1lsc.mdl
lrwxrwxrwx 1 controls controls      54 Apr 28 14:41 c1ioo.mdl -> /opt/rtcds/cds_user_apps/trunk/ISC/c1/models/c1ioo.mdl
lrwxrwxrwx 1 controls controls      54 Apr 28 14:41 c1gpv.mdl -> /opt/rtcds/cds_user_apps/trunk/ISC/c1/models/c1gpv.mdl
lrwxrwxrwx 1 controls controls      54 Apr 28 14:41 c1gfd.mdl -> /opt/rtcds/cds_user_apps/trunk/ISC/c1/models/c1gfd.mdl
lrwxrwxrwx 1 controls controls      54 Apr 28 14:41 c1gcv.mdl -> /opt/rtcds/cds_user_apps/trunk/ISC/c1/models/c1gcv.mdl
lrwxrwxrwx 1 controls controls      54 Apr 28 14:41 c1ass.mdl -> /opt/rtcds/cds_user_apps/trunk/ISC/c1/models/c1ass.mdl

  4582   Thu Apr 28 15:31:36 2011 kiwamuUpdateIOOMC PDH lock : readjustment of demodulation phase

Since Suresh has installed the RF source box and changed the cable configuration somewhat,

the demodulation phase for the MC locking became off by about 10 degree.

I changed the length of some cables and obtained a good demodulation phase by the same technique as Suresh and Koji did before (see here for detail).

I maximized the Q signal. The lock of the MC looks healthy.

DSC_2975_ss.jpg

Quote from #4578

RF Source box has been mounted in the 1X2 rack. 

  4581   Thu Apr 28 12:25:11 2011 josephbUpdateCDSFurther adventures in Hyper-threading

First, I disabled front end starts on boot up, and brought c1sus up.  I rebuilt the models for the c1sus computer so they had a new specific_cpu numbers, making the assumption that 0-1 were one real core, 2-3 were another, etc.

Then I ran the startc1SYS scripts one by one to bring up the models.  Upon just loading the c1x02 on "core 2" (the IOP), I saw it fluctuate from about 5 to 12.  After bringing up c1sus on "core 3", I saw the IOP settle down to about 7 consistently.  Prior to hyper-threading it was generally 5. 

Unfortunately, the c1sus model was between 60 and 70 microseconds, and was producing error messages a few times a second

[ 1052.876368] c1sus: cycle 14432 time 65; adcWait 0; write1 0; write2 0; longest write2 0
[ 1052.936698] c1sus: cycle 15421 time 74; adcWait 0; write1 0; write2 0; longest write2 0

Bringing up the rest of the models (c1mcs on 4, c1rfm on 5, and c1pem on 6), saw c1mcs occasionally jumping above the 60 microsecond line, perhaps once a minute.   It was generally hovering around 45 microseconds.  Prior to hyper-threading it was around 25-28 microseconds.

c1rfm was rock solid at 38, which it was prior to hyper-threading.  This is most likely due to the fact it has almost no calculation and only RFM reads slowing it down.

c1pem continued to use negligible time, 3 microseconds out of its 480.

I tried moving c1sus to core 8 from core 3, which seemed to bring it to the 58 to 65 microsecond range, with long cycles every few seconds.

 

I built 5 dummy models (dua on 7, dub on 9, duc on 10, dud on 11, due on 1) to ensure that each virtual core had a model on it, to see if it helped with stabilizing things.  The models were basically copies of the c1pem model.

Interestingly, c1mcs seemed to get somewhat better and only taking to 30-32 microseconds, although still not as good as its pre-hyper-threading 25-28.  Over the course of several minutes it was no longer having a long cycle.

c1sus got worse again, and was running long cycles 4-5 times a second.

 

At this point, without surgery on which models are controlling which optics (i.e. splitting the c1sus model up) I am not able to have hyper-threading on and have things working.  I am proceeding to revert the control models and c1sus computer to the hyper-threading state.

 

 

  4580   Thu Apr 28 10:53:50 2011 josephbUpdateCDSAdventures in Hyper-threading

What was done:

1) Turn off MC1, MC2, MC3, BS, ITMX, ITMY, PRM, SRM watchdogs.

2) Turn c1sus computer off (sudo shutdown now)

3) Go connect monitor and keyboard to c1sus.  Turn c1sus on.

4) Hit "del" key at the right time to go to setup (BIOS).

5) Go to BIOS advanced tab, CPU options, enable Multi-threading.

6) Hit F10 to save and let the computer continue booting.

What went wrong:

Once c1sus was up, I noticed several red lights and dead keep alives for the c1sus models.

Typing dmesg on c1sus revealed many messages like:

[  107.583420] c1x02: cycle 33737 time 20; adcWait 10; write1 0; write2 0; longest write2 0
[  107.583771] c1x02: cycle 33760 time 19; adcWait 11; write1 0; write2 0; longest write2 0

This indicates the Input/Output Processor (IOP) is not completing its duties within the 15 microseconds (1/64 kHz) that it has.  These lines indicate its take 20 or 19 microseconds.  (I saw messages ranging from 16 to 22 microseconds).

So this seems to agree with Rolf's observations that hyperthreading can cause a 5-10 microsecond increase in computation time.

So the next thing to do is modify which core the codes are running on, and try to get them paired up on the same physical core.

  4579   Thu Apr 28 07:14:34 2011 SureshUpdateRF SystemRF Distribution box installed

RF Distribution box has been mounted in the 1Y2 rack and is ready for use.

 

P4280066.JPG

The box receives 11 and 55 MHz Demod Signals from the RF source located in the 1X2 rack.

  4578   Thu Apr 28 06:46:30 2011 SureshUpdateRF SystemRF Source installed

RF Source box has been mounted in the 1X2 rack. 

P4280064.JPG

 

Heliax cables have been directly attached to the box and anchored on the side of the 1X2 rack.  Here is a list of Helix cables which have been connected so far.

 

Cables old name New name From -> To
1 133 MHz 11 Mhz Demod 1X2 to 1Y2 rack
2 199 MHz 55 MHz Demod 1X2 to 1Y2 rack
3 166 EOM 11 MHz EOM 1X2 to PSL table
4 33 EOM 55 MHz EOM 1X2 to PSL table
5 REFL 33 AS11 AS table to 1Y2

 

  4577   Wed Apr 27 21:19:25 2011 kiwamuUpdateLSCLSC whitening for PD1-4

On the back side of 1Y2 rack I found a cable, CAB-1X2-LSC_7, which is supposed to be connected to the whitening filter was disconnected.

I plugged it back and confirmed that the whitening filter is under control of EPICS.

Now all the gain sliders seem to be working because I can change the amplitude of signals with the sliders.

 

(method)

  To check if the gain sliders are working or not, I intentionally disconnected all the inputs to the whitening filter.

Then I brought a gain slider of interest to the maximum. Due to the big gain I was easily able to see noise lying above ADC noise.

Also if the gain slider is 0 dB, which is the minimum value, the spectrum becomes just ADC noise.

In this way I checked all the gain sliders from PD1 to PD4. The picture below is just an example screenshot when I was doing this test.

Note that each filer is designed to have two poles at 150 Hz and two zeros at 15 Hz.

Screenshot-1.png

Quote from #4570

While checking whitening filters on the LSC rack, I found some epics controls for the whitening looked not working.

So I powered two crates off : the top one and the bottom one on 1Y3 rack.

These crates contain c1iscaux and c1iscaux2. Then powered them on. But it didn't solve the issue.

  4576   Wed Apr 27 21:08:08 2011 ranaUpdateLSCAS11

I worked on AS_11 today. Its ready for its noise / optical gain calibrations. I have left it on Suresh's desk.

AS11.png

This was one of the 24.5 MHz Black Box (Ben Abbott) style RFPDs rescued from LLO. The tunable inductor that was installed was too small to get the frequency down to 11 MHz and so I swapped in one of the shielded, ferrite core ones from our '7mm' CoilCraft kit. It had a range of 1.2 - 1.8 uH according to the datasheet.

I wasn't able to simulataneously get the peak at 11.06 MHz and the notch at 59.3 MHz and so I took Koji's advice and tuned the peak best. The plot above shows how the notch is slightly off. I think its not a problem; to get it better we would have to change out the inductor for the "2-omega" notch, but I was too lazy. The thinking is that its more important to have the gain be symmetric around the signal readout frequency so as to not imbalance the audio sidebands.

Since this one is going to be AS_11, we think that the 22 MHz signal will be tiny: the transmission of the 11 MHz sidebands to the dark port is small. If we later want to put in a 22 MHz notch anyway, there is space to do this via the 'active notch' pads around the MAX4107.

For the above plot, I used the Jenne laser. The DC output of the PD was ~30 mV (~0.6 mA). The RF drive to the laser was -10 dBm: no saturations. I have calibrated out the cable responses, but not using the 1811 setup, so the absolute calibration has yet to be done.

Also, it needs some new stickers. It would be handy if someone can figure out how to get some sheets of stickers that we can put into the printer. Then we can laser printer all of the data onto the stickers and stick them to the RFPD box.

  4575   Wed Apr 27 20:14:16 2011 AidanSummaryelogRestarted with script ...
  4574   Wed Apr 27 18:14:48 2011 kiwamuUpdateLSCpreparation for DRMI locking : RF status

RF_Work_Status.png

POX11 (see this entry) is now listed as REFL11 (on the very top row).

We will rename POY11 to POP11 for DRMI locking.

The files are on https://nodus.ligo.caltech.edu:30889/svn/trunk/suresh/40m_RF_upgrade/.

  4573   Wed Apr 27 17:38:01 2011 kiwamuUpdateElectronicsRe : AS55 demod board with new 90 deg splitter : healthy

relativephase.png

Figure.1  I-Q relative phase measurement as a function of LO power.

 Blue curve : relative phase of AS55 that I have modified today (#4572).

 Red curve : relative phase of AS11 that I had modified a week ago (#4554). Just for comparison.

 The relative phase of AS55 agrees approximately what we expected according to the datasheet of PSCQ-2-51W. We expected 85 degree.

 

IQamplitude.png

Figure.1  I-Q amplitude imbalance as a function of LO power.

From - 5 dBm to 5 dBm in LO power the imbalance is within 3 %.

But the precision of the measurement is also about 2 % (because I used an oscilloscope). Even so the imbalance is still good.

Quote from #4572

Some plots will be posted later.

 

  4572   Wed Apr 27 15:34:38 2011 kiwamuUpdateElectronicsAS55 demod board with new 90 deg splitter : healthy

A new 90 deg splitter, PSCQ-2-51W, has been installed on another demod board called AS55.

It shows a reasonably close 90 degree separation between the I and Q signals at 55 MHz with various LO and RF power.

So far we have ordered only three PSCQ-2-51Ws for test. Now we will order some more for the other demodulators.

 Some plots will be posted later.

  4571   Tue Apr 26 22:56:35 2011 ZachUpdateelogrestarted

 with script

  4570   Tue Apr 26 22:56:01 2011 kiwamuUpdateLSCc1iscaux2 and c1iscaux restrated

While checking whitening filters on the LSC rack, I found some epics controls for the whitening looked not working.

So I powered two crates off : the top one and the bottom one on 1Y3 rack.

These crates contain c1iscaux and c1iscaux2. Then powered them on. But it didn't solve the issue.

  4569   Tue Apr 26 22:03:49 2011 ranaUpdateElectronicsPOX_11 debugging

I used the Jenne AM laser to tune up the PD (used to be POX_11 but now is called REFL_11). In addition to the notch at 22 MHz, I have also put in a LC notch at 5*f = 55.3 MHz. The transfer function below shows the RF OUT of the PD v. the drive to the laser. I didn't divide out by the 1811 because its not on the EE bench.

MM7.png

 

  4568   Tue Apr 26 01:20:02 2011 ranaUpdateElectronicsPOX_11 debugging

The performance plots for POX_11 in the wiki are horrendous and the schematic is missing.

I opened up the box and found all kinds of horrors. There were multiple tunable parts and a flurry of excess nonsense.

The top 2 worst offenders:

1) The main tunable inductor was busted. I removed it and found that the coil was open. Too much indelicate soldering in its vicinity had melted the wire. Someone had put extra inductors and capacitors around it to make it seem as if the PD was working fine, but the noise performance was off by a factor of ~100.

2) The MAX4107 had a 1.4k series resistor. This make the output go through a 1450/50 voltage division which is not nice for the SNR. I removed it.

I then struggled for awhile to get a sensible response. It turned out that the TEST IN input was not giving me a sensible TF. Jenne and I fired the Jenne laser at it and found that the 11 MHz main resonance is there. In the morning I'll finish this off and post more results. I think its going to end up being fine.

We are going to have to take a careful look at all the RFPDs if this one is any indication...

  4567   Mon Apr 25 22:38:49 2011 kiwamuUpdateLSCprepration for DRMI : Y arm flashing
This week is going to be a recycled Michelson week.
As a preparation I did several things today :
 1. Alignment of the Y arm
 2. Alignment of PRM
 3. Checking of all the pick-off ports

 


 
(Y arm alignment)
 The idea to have the Y arm aligned is that : once we lock the Y arm we will be able to align the input PZTs using the Y arm as a reference.
 I tried aligning the Y arm and successfully made the Y arm flashing with IR. I can see it flashing on ITMY camera but no flashing on ETMY camera.
 
(PRM alignment)
PRM has been intentionally misaligned for the single arm green locking test.
I just confirmed that we can bring PRM back to a good alignment. Now we can see the central part is flashing too.
 
(picked-off beams)
I went checking through all the picked off beams to see if they are available or not.
POX : lost
POY : fine
POP : very clipped
POSRM : fine
  4566   Mon Apr 25 12:55:35 2011 AlastairBureaucracyComputerswiki?

Quote:

Quote:

40m wiki seems to have been down for quite a while now but I can't see any info in the elog about it.  Is there some ongoing problem?

 There was an email from Dave Barker about this.  They had to reorganize the DNS at LHO.  The URL that should be used is: http://blue.ligo-wa.caltech.edu:8000/40m

 Thanks Jamie, I've updated the links from the ATF wiki to reflect this.

  4565   Mon Apr 25 12:55:19 2011 JenneBureaucracyComputerswiki?

Quote:

Quote:

40m wiki seems to have been down for quite a while now but I can't see any info in the elog about it.  Is there some ongoing problem?

 There was an email from Dave Barker about this.  They had to reorganize the DNS at LHO.  The URL that should be used is: http://blue.ligo-wa.caltech.edu:8000/40m

 Nope.  I know I had the right address, and it was down for me too all weekend.  It's better now though.  blue.ligo-wa.caltech.edu was up though.

  4564   Mon Apr 25 11:58:37 2011 JamieBureaucracyComputerswiki?

Quote:

40m wiki seems to have been down for quite a while now but I can't see any info in the elog about it.  Is there some ongoing problem?

 There was an email from Dave Barker about this.  They had to reorganize the DNS at LHO.  The URL that should be used is: http://blue.ligo-wa.caltech.edu:8000/40m

  4563   Mon Apr 25 11:23:41 2011 AlastairBureaucracyComputerswiki?

40m wiki seems to have been down for quite a while now but I can't see any info in the elog about it.  Is there some ongoing problem?

119427.strip.gif

  4562   Sun Apr 24 21:37:40 2011 kiwamuUpdateIOOreview of triple resonant EOM : model looks fine

To design a new resonant EOM box I started reviewing the prototype that I've built.

As a part of reviewing I checked an important thing that I haven't carefully done so far :

I compared the measured input impedance with that of predicted from a circuit model. I found that they show a good agreement.

So I am now confident that we can predict / design a new circuit performance.


* * * (input impedance) * * *

 Performance of a resonant circuit is close related to its input impedance and hence, in other words, determined by the input impedance.

Therefore an investigation of input impedance is a way to check the performance of a circuit. That's why I always use impedance for checking the circuit.

The plot below is a comparison of input impedance for the measured one and one predicted from a model. They show a good agreement.

(Note that the input impedance is supposed to have 50 Ohm peaks at 11, 29.5 and 55 MHz.)
Input_impedance.png

 
* * * (circuit model) * * *

To make the things simpler I assume the following three conditions in my model:

 1. inductor's loss is dominated by its DC resistance (DCR)

 2. capacitor's loss is characterized only by Q-value

 3. Transformer's loss is dominated by DCR and its leakage inductance

All the parameters are quoted from either datasheet or my measurement. The model I am using is depicted in the schematic below.

Basically the Q-vaules for the capacitors that I used are quite low. I think higher Q capacitors will improve the performance and bring them to more 50 Ohm.

EOMcircuitmodel.png
 

  4561   Fri Apr 22 12:07:38 2011 josephb, steveUpdateCDSRemoved hanging D-sub to SCSI in 1X2

Problem:

Way back, Jay had D-sub to SCSI adapters made to adapt our existing Sander box AA filters to the new SCSI based IO chassis.  However, these did not fit inside the box.

At the time, we simply left the cards outside hanging, which was a hack and needed to be replaced.

Solution:

Steve modified a black AA filter box so that it could fit the D-sub to SCSI adapter board on it, plus strain relief the SCSI cable, rather than let it hang.  The back of the box was cut, and an extending piece of metal attached to the bottom of the box.  The adapter board was screwed into the box, the SCSI plugged in, then the SCSI cable is clamped to the extending metal as well.

This modification will be propagated to the 3 remaining AA filter boards using the D-sub to SCSI adapter.

  4560   Fri Apr 22 11:08:50 2011 kiwamuUpdateLSCdemod board AS11 : amplitude imbalance

Amplitude imbalance between I and Q in a demod board, AS11, with the new 90 deg splitter was measured.

It shows roughly 10% amplitude imbalance when the LO power is in a range from 0 to 5 dBm. Not so bad.

 

  With the handmade coil there used to be a huge imbalance (either I or Q goes to zero volt while the other keeps about 1 V rms) as the LO power decreases.

But with the new 90 deg splitter now there are no more such a huge imbalance.

The remaining 10 % imbalance possibly comes from the fact that we are using ERA-5 in each I and Q path. They may have such gain imbalance of 10%.

We should check the ERA-5 gains so that we can confidently say ERA-5 causes the amplitude imbalance.

Then our plan replacing the ERA-5s (see here) will sound more reasonable.

IQamplitude.png

 

Quote from #4555

The new 90 deg splitter works better.

 I will also measure amplitude unbalances between I and Q.

 

  4559   Fri Apr 22 10:28:22 2011 ranaUpdateRF SystemRF Source Harmonics
You should be able to resolve the other harmonics by decreasing the IF BW or RBW on the analyzer. Even though
they're OK, its useful to have the final measurement of all of them in some kinds of physical units (like dBm, but
not dBm/Hz or dB or dBcubits).
  4558   Fri Apr 22 09:25:43 2011 SureshUpdateRF SystemRF Source: Temperature sensor relocated

RF Amp operating temperature

Earlier measurement reported by Alberto in LIGO-T10004-61-v1 based on the LM34 temperature sensor were lower than that shown by placing a calibrated thermocouple sensor directly on the heat sink by about 5deg C. The difference probably arose because the LM34 was located on a separate free-hanging copper sheet attached to the RF Amp by a single screw, resulting in a gradient across the copper strip.   I tried to move the LM34 which was glued down, but broke the leads in the process.  I then replaced it with another one mounted much closer to the heat sink and held it down with a copper-strip clamp.  There is no glue involved and there is heatsink compound between the flat surface of the LM34 and the heatsink.  Picture attached. 

  The picture also shows the new filters which have been put in place to reduce the harmonics.  Note that the SBP-10.7 which was to go on the 11 MHz Demod output is located much farther upsteam due to space constraints.

P4220056.JPG

  4557   Fri Apr 22 09:05:53 2011 SureshUpdateRF SystemRF Source Harmonics
As seen in the previous measurement the first harmonic of both the 11 MHz and 55 MHz outputs are about 30dB
higher than desired.  In an attempt to attenuate these and higher harmonics I introduced SBP-10.7 filters into
the 11MHz outputs and SLP-50 filters into the 55 MHz outputs.
Then I measured the height of the harmonics again and found that they were suppressed as expected.  Now harmonic
at 22 MHz is 58dB lower than the 11 MHz fundamental.  And the 110 MHz is lower by 55 dB compared to the 55 MHz
fundamental.  None of the higher harmonics are seen => they are below 70dB

SLP-50 has an insertion loss(IL) of 4.65 dB and Return Loss(RL) of 3dB.  It would be better to use SBP-60
(IL=1.4 dB and RL=23dB)

The filter on the 11 MHz lines is okay. The SBP-10.7 has IL=0.6 dB and RL=23 dB.
  4556   Fri Apr 22 02:10:53 2011 ZachUpdateelogrestarted

Restarted the elog with the script.

  4555   Thu Apr 21 21:46:22 2011 kiwamuUpdateLSCdemod board : new 90 deg splitter

A less LO power dependence on the relative phase was found. The new 90 deg splitter works better.

From -3 dBm to 10 dBm in LO power, the relative phase is within 90 +/- 5 deg.

As a comparison I plot the phase that I measured when the handmade coil had been there (green curve in the plot).

relativephase.png

 

 I will also measure amplitude unbalances between I and Q.

Quote from #4554

A 90 degree splitter, PSCQ-2-51W, has arrived today and I installed it on a demod board called AS11.

Results of the I-Q phase measurement with the new splitter will be reported soon.

 

 

  4554   Thu Apr 21 21:24:41 2011 kiwamuUpdateLSCdemod board : new 90 deg splitter

A new 90 degree splitter, PSCQ-2-51W, has arrived today and I installed it on a demod board called AS11.

Results of the I-Q phase measurement with the new splitter will be reported soon.

 

 * Picture 1 = before removal of the handmade coil

 * Picture 2 = after removal of the coil and the associated capacitors

 * Picture 3 = after soldering PSCQ-2-51-W

DSC_2949_ss.jpg

DSC_2951.JPG

DSC_2952_ss.jpg

Quote from #4358

 First of all we will replace the home-made 90 degree splitter (see this entry) by a commercial splitter, PSCQ-2-51-W+ from Mini circuit. This is the step 1 basically.

  4553   Thu Apr 21 15:20:46 2011 steveUpdatePEMPSL enclosure work today

Quote:

The east side window guides  will be replaced by one long U-channel. There will be drilling into 2x2 steel frame. It should be done by 2pm today

This should remove the jerking motion of windows hitting the individual guides.

 The installation went smoothly. There will be no more banging the doors into  guides on the east side.

Atm2 showing the west side as is today

Attachment 1: P1070558.JPG
P1070558.JPG
Attachment 2: P1070560.JPG
P1070560.JPG
  4552   Thu Apr 21 15:03:29 2011 steveUpdateComputersjuction board finds home

The anti aliasing box was opened up at the back to accommodate the junction board and the SCSI cable towards the ADC. Aluminum plate was attached to the bottom to hold the strain relief clamp.

Three more hanging junction cards will be replaced in this manner.

Attachment 1: P1070570.JPG
P1070570.JPG
Attachment 2: P1070568.JPG
P1070568.JPG
  4551   Thu Apr 21 14:39:43 2011 steveUpdateRF Systemnew strain relieved N connectors at AP

New right angle PVC, 2 x 2 x  1/4" installed at the AP table to strain relief the 1/4" spiral corrugated RF coaxes.

Attachment 1: P1070562.JPG
P1070562.JPG
Attachment 2: P1070564.JPG
P1070564.JPG
  4550   Thu Apr 21 08:26:42 2011 steveUpdatePEMPSL enclosure work today

The east side window guides  will be replaced by one long U-channel. There will be drilling into 2x2 steel frame. It should be done by 2pm today

This should remove the jerking motion of windows hitting the individual guides.

  4549   Wed Apr 20 23:20:49 2011 jamieSummaryComputersinstallation of CDS tools on pianosa

This is an overview of how I got (almost) all the CDS tools running on pianosa, the new Ubuntu 10.04 control room work station.

This is machine is experiment in minimizing the amount of custom configuration and source code compiling. I am attempting to install as many tools as possible from existing packages in

available packages

I was able to install a number of packages directly from the ubuntu archives, including fftw, grace, and ROOT:

apt-get install \
libfftw3-dev \
grace \
root-system

LSCSOFT

I installed all needed LSCSOFT packages (framecpp, libframe, metaio) from the well-maintained UWM LSCSOFT repository.

$ cat /etc/apt/sources.list.d/lscsoft.list
deb http://www.lsc-group.phys.uwm.edu/daswg/download/software/debian/ squeeze
deb-src http://www.lsc-group.phys.uwm.edu/daswg/download/software/debian/ squeeze contrib
sudo apt-get install lscsoft-archive-keyring
sudo apt-get update
sudo apt-get install ldas-tools-framecpp-dev libframe-dev libmetaio-dev lscsoft-user-en

You then need to source /opt/lscsoft/lscsoft-user-env.sh to use these packages.

EPICS

There actually appear to be a couple of projects that are trying to provide debs of EPICS. I was able to actually get epics working from one of them, but it didn't include some of the other needed packages (such as MEDM and BURT) so I fell back to using Keith's pre-build binary tarball.

Prereqs:

apt-get install \
libmotif-dev \
libxt-dev \
libxmu-dev \
libxprintutil-dev \
libxpm-dev \
libz-dev \
libxaw7-dev \
libpng-dev \
libgd2-xpm-dev \
libbz2-dev \
libssl-dev \
liblapack-dev \
gfortran

Pulled Keith's prebuild binary:

cd /ligo/apps
wget https://llocds.ligo-la.caltech.edu/daq/software/binary/apps/ubuntu/epics-3.14.10-ubuntu.tar.gz
tar zxf epics-3.14.10-ubuntu.tar.gz

GDS

I built GDS from svn, after I fixed some broken stuff [0]:

cd ~controls/src/gds
svn co https://redoubt.ligo-wa.caltech.edu/svn/gds/trunk
cd trunk
#fixed broken stuff [0]
source /opt/lscsoft/lscsoft-user-env.sh
./bootstrap
export GDSBUILD=online
export ROOTSYS=/usr
./configure --prefix=/ligo/apps/gds --enable-only-dtt --with-epics=/ligo/apps/epics-3.14.10
make
make install

dataviewer

I installed dataviewer from source:

cd ~controls/src/advLigoRTS
svn co https://redoubt.ligo-wa.caltech.edu/svn/advLigoRTS/trunk
cd trunk/src/dv
#fix stupid makefile /opt/rtapps --> /ligo/apps
make
make install

I found that the actual dataviewer wrapper script was also broken, so I made a new one:

$ cat /ligo/apps/dv/dataviewer
#!/bin/bash
export DVPATH=/ligo/apps/dv
ID=$$
DCDIR=/tmp/${ID}DC
mkdir $DCDIR
trap "rm -rf $DCDIR" EXIT
$DVPATH/dc3 -s ${NDSSERVER} -a $ID -b $DVPATH "$@"

environment

Finally, I made a environment definer file:

$ cat /ligo/apps/cds-user-env.sh
# source the lscsoft environment
. /opt/lscsoft/lscsoft-user-env.sh

# source the gds environment
. /ligo/apps/gds/etc/gds-user-env.sh

# special local epics setup
EPICS=/ligo/apps/epics
export LD_LIBRARY_PATH=${EPICS}/base/lib/linux-x86_64:$LD_LIBRARY_PATH
export LD_LIBRARY_PATH=${EPICS}/extensions/lib/linux-x86_64:$LD_LIBRARY_PATH
export LD_LIBRARY_PATH=${EPICS}/modules/seq/lib/linux-x86_64:$LD_LIBRARY_PATH
export PATH=${EPICS}/base/bin/linux-x86_64:$PATH
export PATH=${EPICS}/extensions/bin/linux-x86_64:$PATH
export PATH=${EPICS}/modules/seq/bin/linux-x86_64:$PATH

# dataviewer path
export PATH=/ligo/apps/dv:${PATH}

# specify the NDS server
export NDSSERVER=fb

[0] GDS was not compiling, because of what looked like bugs. I'm not sure why I'm the first person to catch these things. Stricter compiler?

To fix the following compile error:

TLGExport.cc:1337: error: ‘atoi’ was not declared in this scope

I made the following patch:

Index: /home/controls/src/gds/trunk/GUI/dttview/TLGExport.cc
===================================================================
--- /home/controls/src/gds/trunk/GUI/dttview/TLGExport.cc (revision 6423)
+++ /home/controls/src/gds/trunk/GUI/dttview/TLGExport.cc (working copy)
@@ -31,6 +31,7 @@
#include <iomanip>

#include <string.h>

#include <strings.h>

+#include <stdlib.h>


namespace ligogui {
using namespace std;

To fix the following compile error:

TLGPrint.cc:264: error: call of overloaded ‘abs(Int_t&)’ is ambiguous

I made the following patch:

Index: /home/controls/src/gds/trunk/GUI/dttview/TLGPrint.cc
===================================================================
--- /home/controls/src/gds/trunk/GUI/dttview/TLGPrint.cc (revision 6423)
+++ /home/controls/src/gds/trunk/GUI/dttview/TLGPrint.cc (working copy)
@@ -22,6 +22,7 @@
#include <fstream>

#include <map>
#include <cmath>

+#include <cstdlib>


namespace ligogui {
using namespace std;

  4548   Wed Apr 20 22:29:07 2011 sureshUpdateRF SystemPlan for LSC rack

The suggested layout of the 1Y2 Rack is shown below.

To simplify the wiring, I have largely kept demod boards with the same same LO frequency close to each other. 

The Heliax cables land on the top and bottom of the of subracks.  These are currently flexible plastic sheets.  Steve has agreed to replace them with something more rigid.  It would be good to have eight N-type connectors on the top and eight  at the bottom.  As  demod boards occur in sets of eight per subrack.  So it would be convenient if the 11 and 55 Mhz Heliax cables land on the top and the rest at the bottom.  In the layout I have shown the current situation. 

The LO signals to the boards come from the RF Distribution box and this is kept in the middle so that cables to both the subracks can be kept short.

The outputs of the AA filter boards from both subracks  have to be connected to the SCSI Interface board with a twisted pair ribbon cable. 

1Y2_Rack_Layout.png

  4547   Wed Apr 20 21:53:01 2011 SureshConfigurationRF SystemRF system: Stray heliax cable

We found a stray unused heliax cable running from the LSC rack 1Y2 to a point between the cabinets 1X3 and 1X4. This cable will need to be redirected to the AS table in the new scheme.   It is labled C1LSC-PD5  The current situation has been updated as seen in the layout below

rogue_cable_1.png

Attachment 1: rogue_cable_1.png
rogue_cable_1.png
  4546   Wed Apr 20 20:42:55 2011 kiwamuUpdateVIDEOtoday's video session

[Steve / Suresh / Kiwamu]

90 % of unused video cables have been removed.

Still a couple of video cables are floating around the video MUX. They will be removed in the next week's session.

DSC_2931_ss.jpg

DSC_2938_ss.jpg

 

ELOG V3.1.3-